In the area around Iqra University, a number of food vendors gather each lunchtime to sell food

to university students who are tired of Hostel food. The university and the city have no license

fees that apply to food vendors, preferring to let the market dictate how many and which vendors

show up.

Many different cuisines are represented on the street corner, including a cart sponsored by Tipu

Burger sells a burger of either Chicken or Beef for Rs.100. The food is made in the morning at

nearby restaurant in DHA, when the kitchen is otherwise unoccupied.

His crew of three, each of whom earns Rs.200 per hour, takes 2 hours to make the 50 meals

required by Tipu. In creating these burgers, they use ingredients equal to Rs.2000. Tipu hires

another worker to load her cart with food and sell it during the lunch hours. That worker costs

Rs.100 per hour and typically sells out the entire cart of 50 meals in 2 hours. The cart is rented

for Rs.1000 per 5-day week. (The carts are not in operation on the weekends, when Tipu is too

busy at his restaurant.)

A. What market structure does this business most resemble? What characteristics lead you to

this conclusion? (3 Marks)

B. What would you expect to see happen in this business? Use the data in the problem to

support your conclusions. (2 Marks)


1
Expert's answer
2021-01-27T18:46:25-0500


a) The market structure being represented in the business has resemblance to perfect competition. This is because there are many buyers and sellers. There is no restriction on entry or exit by firms.




b) New vendors will enter the business as there is positive profit and profit will fall as new vendors enter.



b) Currently Tipu's costs (5-day week) = 5(200*2*3+2000+100*2)+1000=18000


His revenue (5-day week)=5(50*100)=25000


so, his profit=25000-18000=7000




As the businesses are making positive profit, and the market structure is that of perfect competition, it will attract new vendors until the profits are brought down to zero which is the long-run equilibrium in perfectly competitive markets.